9. Use advancement operators to find a closed-form formula for xn, as a 0, function of n, given that forn2 and that

applied combinatorics
Answer #1

Given reccurence relation is

x_{n+3}=7x_{n+2}-16x_{n+1}+12x_n+2^n___(*)

Since the given reccurence relation is non homogenuous ,

Guess

y_n=c2^n+d\\ y_{n+3}=c2^{n+3}+d\\ y_{n+2}=c2^{n+2}+d\\ y_{n+1}=c2^{n+1}+d\\

Now put all the values in (*).

n +3

                        =7c2^{n+2}+7d-16(c2^{n+1}-16d+12c2^{n}+12d+2^n

                         =7c2^{n+2}-16c2^{n+1}+12c2^{n}+2^n+3d

7c2^{n+2}-16c2^{n+1}+12c2^{n}+2^n-c2^{n+3}+2d=0

7c2^{n+2}-16c2^{n+1}+(1+12c)2^{n}-c2^{n+3}+2d=0

Now compare the coef ,

1+12c=0,d=0\\ c=-1/12,d=0

Hence

1+12c=0,d=0\\ c=-1/12,d=0\\ y_n=-1/12(2)^nis the particular soolution .

For homogenuous ,

h_{n+3}=7h_{n+2}-16h_{n+1}+12h_n

1=7r-16r^2+12r^3

(2r-1 )2(3r-1)-0

h_n=(A+Bn)(1/2)^n+C(1/3)^n

Hence solution is

x_n=y_n+h_n=-1/12(2)^n+(A+Bn)(1/2)^n+C(1/3)^n

Intial conditions

x_0=0=-1/12(2)^0+(A+B*0)(1/2)^0+C(1/3)^0

                    =-1/12+A+C\\ A+C=1/12--(1)

x_1=0=-1/12(2)^1+(A+B*1)(1/2)^1+C(1/3)^1

                    xi = 0 =-1/6 + (A + B)/2 + C/3

x_2=1=-1/12(2)^2+(A+B*2)(1/2)^2+C(1/3)^2

                  1=-1/3+(A+2B)/4+C/9

(9A+18B)+4C=12--(3)

By (1),(2),(3)

A=-20/3,B=5/2,C=27/4

Hence